OpenStudy (faman39):

Please anyone in here, who love math and good at it, i will give them who able to answer correctly all of my math problem with clear explanation, i will give medal, who will able help me with all like that. Please need to solve my all math problems and hav e too understand, if i wont to pass my upcoming exams. So, please help me, i will give medals for all. a. m - (-5) =13. b. v - 9 =4. c. -5+s =13. d. g + (-4) = 9.

OpenStudy (karatechopper):

lets start with m-(-5)=13 well we can add-the-opposite! and make that m+ (+5)=13 so then it would obviously be m+5=13 now to solve you subtract 5 from both sides leaving you with m=8

OpenStudy (karatechopper):

understood?

OpenStudy (anonymous):

\[M= 13 + 5 = 1\]A. The reason you do this is because you are pretty much subtracting 5 thats it and all your gonna have to do is add 13 and 5 so that you dont subtract wrong... SO m= 18

OpenStudy (anonymous):

A. 2 negatives equal a positive so m plus 5=13 then subtract 5 m+8

OpenStudy (anonymous):

Then B\[V = 9 + 4 so V= 13\] the reason this is is because you add the two numbers to get a higher number subtracted from that gives you the answer.

OpenStudy (anonymous):

m=8 sorry not plus

OpenStudy (anonymous):

\[-5 + 5 = 0 0-13 = 13 so 13 +5 = 18 \S= 18\] C

OpenStudy (anonymous):

So S = 18

OpenStudy (faman39):

thank u so much guys for help! i will give medals for sure!

OpenStudy (anonymous):

And to d.\[so 9 + -4 is -4 - 0 so 4 then 9+4 = 13 \] So G=13

OpenStudy (anonymous):

Hope that helps:)